# MMRA: A Benchmark for Multi-granularity Multi-image Relational Association

### v1

@@ Abstract
Given the remarkable success that large visual language models (LVLMs) have achieved in image perception tasks, the endeavor to make LVMLs perceive the world like humans is drawing increasing attention. Current multi-modal benchmarks mainly focus on the objective fact or certain topic related potential knowledge within a image, but overlook the associative relations between multiple images. Therefore, we define a multi-image relation association task, and meticulously curate MMRA benchmark, a Multi-granularity Multi-image Relational Association benchmark, consisted of 1026 samples. In order to systematically and comprehensively evaluate mainstream LVLMs, we establish an associational relation system among images that contain 11 subtasks (e.g, UsageSimilarity, SubEvent, etc.) at two granularity levels (i.e., ”image” and ”entity”) according to the relations in ConceptNet. Our experiments demonstrate that, on our MMRA benchmark, current mainstream LVLMs all have their own advantages and disadvantages across different subtasks. It is worth noting that, at the entity level, the performance of all models is worse than that of them at the image level, indicating that the fine-grained multi-image perception task is still challenging for LVLMs. The tasks related to spatial perception are relatively difficult for LVLMs to handle. Furthermore, we find that LVMLs exhibit a good ability to perceive image details, and the key to enhancing their multi-image association capability is to strengthen the reasoning ability of their language model component. All our codes and data are released at htthttps://github.com/Wusiwei0410/MMRA.
@@
Figure 1: Overview of the MMRA benchmark. The left part shows the relations extended from the ConceptNet. The right part presents the relation mining process under our Relation System.
@@ 1 Introduction
Multi-modal perception is a crucial factor for achieving general Artificial Intelligence (AI) that can perceive the world similarly to humans. The multi-modal perception has made significant advancements due to the development of Large Visual Language Models (LVLMs). Recent multi-modal LVLMs have improved the processing of visual information through the use of diverse visual encoders and various alignment techniques to bridge the gap between text and image modalities. Consequently, there is growing interest in systematically and comprehensively defining a benchmark to assess the performance of LVLMs and guide future development in this field. The current multi-modal benchmarks focus on asking about objective facts within a single image (i.e., the number of gloves in the picture) or the underlying causes of events that require reasoning to ascertain (i.e., the reason the grass on the prairie is moving is the wind). However, those benchmarks all overlooked the evaluation of the multi-image association ability of LVMLs (e.g., ’those images all depict outdoor scenes’ as shown in Fig 1).
This ability to identify possible connections between various elements in the world is crucial for human perception. To explore the multi-image perception capabilities of LVMLs, we have defined a multi-image relation association task. This task requires LVLMs to discern the potential relations between two images (for instance, recognizing that both the knife in Image 1 and the car bodies in Image 2 are made of iron).
@@
Exploring the potential relations between events has received widespread attention in the field of natural language processing (NLP) [5, 31, 7, 12]. This association-based commonsense reasoning ability is widely used to measure the capability of LLMs (Large Language Models) [4]. However, there is no comprehensive multi-image benchmark systematically defining the association relations among multiple images to guide the development of multi-image models. On the one hand, current multi-image benchmarks [30, 26] only focus on monotonous forms of tasks which ask question about objective facts among pictures like VQA task (e.g., how many gloves are there in all pictures), while neglecting the potential connections between the images (for example, that the material composing the entities in the images is wood). On the other hand, the difficulty of mining relations among multiple images at different granularities (i.e., entity level and image level) varies differences, most tasks in those benchmarks all focus on the image level which require inference through the overall content of the images (e.g., whether two images depict consecutive events in the same scene). Although the VCD systematically define the potential knowledge of entities in a image, it just focus on the relation between the entity and the commonsense knowledge and cannot directly be applied to define the relation among images.
@@
To address the aforementioned challenges, we manually curate the MMRA benchmark, a Multi-granularity Multi-image Relational Association benchmark. Based on the relations in ConceptNet and observations of potential connections between images, we have defined a associational relation system, which consists of 6 tasks at the entity level (RelativePosition, NearSameEntity, MentalitySimilarity, AppearanceSimilarity, SimilarMaterial and UsageSimilarity) and 5 tasks at the image level (Layout, Environments, SimilarEvent, SubEvent and ObservationAngle) across from different perspectives of mining relation between images (outline in Fig 1). Specifically, we select a part of the images in LLaVA-665k-multi dataset and employ 5 postgraduate students to manually label 1,026 image pairs with questions and answers on the selected data. Besides, we also conduct the elimination of answer leakage and prove the necessity of it.
We evaluate LVMLs and LLMs using our MMRA benchmark across four different input settings: Image+Question (IQ), Description+Question (DQ), Image+Description+Question (IDQ), and Question Only (QO). We present our key insights as follows:
@@
    Based on the results of the IQ and QO setting, we found that closed-source models like GPT-4o, GPT-4v and Gemini-Flash outperformed all open-source models. In particular, GPT-4o achieved state-of-the-art (SOTA) performance in the overall evaluation. Additionally, different models exhibited significant performance variations across different subtasks of our benchmark. Some open-source models even surpassed GPT-4 in certain subtasks.
@@
    It is worth mentioning that models generally performed better on image level tasks than on entity level tasks and the results on the tasks relevant to spatial perception. It indicates that current LVLMs have weak fine-grained multi-image association capabilities and are not proficient in handling spatial perception tasks.
@@
    As for image-level tasks, after adding image descriptions, the performance of LLMs increases dramatically, second only to GPT-4o and GPT-4v, while the performance of LVLMs does not show any significant change. This indicates that the image-level capability of LVLMs mainly relies on the image content perception ability and the limitation of current LVMLs if the reasoning ability of the language model component. On contrast, after adding image descriptions, the performance of VLM showed significant improvement only in the MS task among all entity-level tasks. This indicates that establishing an emotional connection with the characters in the images requires a good integration of both the image module and the text module.
These findings demonstrate the inadequacies of current LVMLs in multi-image understanding and underscore that the multi-image perceptual demands at the entity and highlight that the requirements for the visual module and the textual module differ when it comes to multi-image association at the entity and image levels.
@@ 2.1 Image Pair Selection
Because most tasks in MMRA benchmark require a certain relation between the given pairs of images, we use the semantic similarity of the images’ caption to select the relatively higher related image pairs for the sake of reducing the difficulty of annotation. To be specific, we randomly chose the images in the LLaVA-665k-multi to form an image pair. We then utilize the SentenceBERT[19] to calculate the semantic similarity to filter the image pair with a score below 0.5. Finally, we obtain 3,403 images pair for annotation.
@@ 2.2 Subtask Definition
Based on the perspective of humans observing images, we divide our tasks into two granularity levels (i.e., entity and the whole image). Most of our subtasks are extended from the relations in the ConceptNet. Additionally, we have also designed some subtasks from a visual perspective (i.e., Layout and ObservationAngle).
As for the entity level.
@@
  • NearSameEntity (NSE): The relation ’LocatedNear’ in ConceptNet expresses ”A and B are typically found near each other”. Based on it, we design a subtask, ’NearSameEntity’, which requires LVLMs to determine whether there are entities, respectively in different images, near the same object.
  • MentalitySimilarity (MS): ’HasProperty’ in ConceptNet is a relation that describes the characteristic of an entity. We think the emotional property expressed by the images could directly affect humans. Thus, we extend this relation to a subtask that requires LVLMs to determine whether the creatures in two images have similar emotions, attitudes, or feelings (e.g., happy, excited, serious, surprised, etc.).
@@
  • SimilarMaterial (SM): The relation ’MadeOf’ expresses ’A is made of B’. Therefore, we design the subtask ’SimilarMaterial’ which requires LVLMs to judge whether there are entities, respectively in different images, with the same production materials.
@@
  • Layout (LO): At the image granularity, we regard the layout of the image as a representation of the relation ”AtLocation”. We design a subtask that requires the LVLMs to determine whether there are similarities in layout between images according to the relation ’NearBy’.
  • Environment (Env): From the visual perspective, the environment of the image is also an important content that humans tend to notice (e.g., both images depict the streets of a European country with a Gothic architectural style). So, we design a subtask lets LVLMs judge if the environments in those images are similar according to the relation ’AtLocation.
  • SimilarEvent (SimE): Excepting the ’SubEvent’ relation in ConceptNet, the similar event is also an crucial factor. So we devise a subtask to evaluate the LVLMs’ capability to find the same event that happened in the given two images.
  • SubEvent (SubE): The temporary relation is an important connection between two images. Therefore, we extend the relation ’SubEvent’ to a subtask requires LVLMs to determine whether the two images describe events that occurred at the same scene in two consecutive moments.
@@ 2.3 Data Annotation
Each student is provided with two images and a specific task. Their responsibility is to determine whether they could design a question based on the given task for the image pair. If the image pair meets the task requirements, they proceed to annotate a question, answer, and options (either multiple-choice or true/false) for that pair. Each annotator cease annotating a task once they reach a predetermined number of labeled samples (i.e., 90) or once all the image pairs for that task have been annotated.
@@
We also conducted cross-validation on the annotated data. Specifically, each annotator review 2-3 tasks labeled by their peers. If any annotated samples don’t meet the task requirements or if the answers derived from the images and options don’t match the correct answer, those samples are removed. Additionally, we ensure that the answers to all questions could not be directly inferred from the text alone (i.e., the question and options). Quality control is concluded once all annotators agreed that their verified portion satisfied the specified requirements.
@@
As shown in the Fig 2, we obtain a total of 1,026 annotated samples. To ensure the validity of the benchmark evaluation, we have endeavored to maintain that the number of samples for all tasks is around 90. The ObservationAngle task has the highest proportion in the entire benchmark, with a total of 126 samples (12.28%). Due to the difficulty of labeling in the NearBy task, we removed some samples with inconsistent opinions from different annotators during the quality control process.
@@ 2.4 Elimination of answer leakage from questions and options
When designing multiple-choice options at the entity level, we need to identify potential entities that could be regarded as the correct answer of question and provide justifications. For example, as illustrated in Fig 1, ’both tree and bench are made of wood’ can be the answer of the SimilarMaterial subtask. However, language models can sometimes deduce the correct answer simply by analyzing the textual content in options. Additionally, annotators often unconsciously label the correct answer with greater detail and specificity, and the language model towards choosing these more detailed options.
@@
Figure 3: Comparison Chart of Results Before and After Text Answer Leak Elimination.
@@ 3.1 Experiments Setting
To explore the impact of LVLM’s image-captioning ability on its multi-image perception, we design four inputs setting: Image + Question, Description + Question, Image + Description + Question and Question Only. Specifically, we use the LLaVA-NeXT-100B to obtain the detailed textual description of the image in the MMRA benchmark.
@@
For the sake of inspecting whether the answer of the questions in our benchmark is leaked in the textual information of options and question, we only input the question to let LVLMs answer.
@@ 3.2 Baselines
As show in Tab 1, we evaluated our benchmark on both mainstream close-source and open-source large models. Regarding close-source LVLMs, we choose the OpenAI’s GPT4o and GPT4v, as well as google’s Gemini-Flash. As for the open-source LVLMs, we mainly evaluate those supporting multi-image inputs (i.e., Idefics2 , Qwen-VL-Chat, Phi3v, Mantis-Idefics2). Besides, we also assess the open-source LLMs (i.e., LLaMA, Qwen1.5-Chat, Qwen2-Chat) under the textual input setting.
@@ 3.3 Evaluation Protocol
As for each task, we all design a prompt to make LVLMs directly generate textual format answers of the question. Except including the content described in the section 3.1, we let LVLMs generate the ’A’, ’B’, ’C’ or ’D’ for the choice questions, and ’Yes’ or ’No’ for the T/F questions. Besides, we also add the options to the prompt. As for further details about our prompt design, please refer to the Tab 4 in Appendix B.
@@
Because the golden answer in our benchmark is in the format of option id (i.e., ’A’, ’B’, ’C’ and ’D’) or judgment (i.e., ’Yes’ or ’No’), we design a rule to match the response of LVLMs with the golden answer. Finally, we use the accuracy of the matching results as the score of those models. Please refer to the Appendix A for the detail of our designed matching rule.

### v2

# MMRA: A Benchmark for Evaluating Multi-Granularity and Multi-Image Relational Association Capabilities in Large Visual Language Models

@@ Abstract
Given the remarkable success that large visual language models (LVLMs) have achieved in image perception tasks, the endeavor to make LVLMs perceive the world like humans is drawing increasing attention. Current multi-modal benchmarks primarily focus on facts or specific topic-related knowledge contained within individual images. However, they often overlook the associative relations between multiple images, which require the identification and analysis of similarities among entities or content present in different images. Therefore, we propose the multi-image relation association task and a meticulously curated Multi-granularity Multi-image Relational Association (MMRA) benchmark, comprising 1,024 samples. In order to systematically and comprehensively evaluate current LVLMs, we establish an associational relation system among images that contain 11 subtasks (e.g, UsageSimilarity, SubEvent, etc.) at two granularity levels (i.e., “image” and “entity”) according to the relations in ConceptNet. Our experiments reveal that on the MMRA benchmark, current multi-image LVLMs exhibit distinct advantages and disadvantages across various subtasks. Notably, fine-grained, entity-level multi-image perception tasks pose a greater challenge for LVLMs compared to image-level tasks. Moreover, LVLMs perform poorly on spatial-related tasks, indicating that LVLMs still have limited spatial awareness. Additionally, our findings indicate that while LVLMs demonstrate a strong capability to perceive image details, enhancing their ability to associate information across multiple images hinges on improving the reasoning capabilities of their language model component. Moreover, we explored the ability of LVLMs to perceive image sequences within the context of our multi-image association task. Our experiments show that the majority of current LVLMs do not adequately model image sequences during the pre-training process. All our codes and data are released at https://github.com/Wusiwei0410/MMRA.
@@
Figure 1: Overview of the MMRA benchmark. Left: image Associational Relations extended from the ConceptNet; Right: the examples of Multi-Image Relation Association task.
@@ 1 Introduction
Multi-modal perception is a crucial factor for achieving Artificial General Intelligence (AGI) that can perceive the world similarly to humans. Significant advancements have been made in multi-modal perception due to the development of Large Visual Language Models (LVLMs) [13, 16, 15, 3, 1], which have improved visual information processing through diverse visual encoders and bridging the gap between text and image modalities by alignment techniques. Consequently, there is growing interest in systematically and comprehensively defining a benchmark to assess the performance of LVLMs and guide future development in this field. The capabilities of LVLMs in associating multi-image relations can more intuitively and systematically reveal potential shortcomings in VLMs when it comes to multi-image perception tasks. (i.e., if LVLMs struggle to determine the spatial relations within images, they are likely to encounter difficulties in accurately perceiving spatial information when interpreting multiple images). However, the current multi-modal benchmarks[24, 30, 31] focus on asking questions within a single image, while the evaluation of LVLMs’ multi-image association ability (e.g., “those images all depict outdoor scenes” as shown in Fig 1) is overlooked.
@@
There is no comprehensive multi-image benchmark systematically defining the association relations among multiple images to guide the development of multi-image models. On the one hand, current multi-image benchmarks [33, 27, 29] only focus on monotonous forms of tasks which ask questions about objective facts among pictures like TextVQA [24] (e.g., how many gloves are there in all pictures), or retrieve similar images in a specific domain such as SciMMIR [29]. All of these benchmarks overlook the potential connections between images (for example, both images contain objects that are made of wood), which could directly reveal the shortcomings of LVLMs from different perspectives. On the other hand, mining relations among multiple images at different granularities (i.e., entity level and image level) and across different topics (e.g., the spatial relation and temporal relation between images) present varying degrees of difficulty, and current multi-image benchmarks have not specifically categorized the tasks based on these distinctions. Although, in the field of Natural Language Processing (NLP), numerous similar works explore the potential relations between textual events or entities [14, 4, 34, 6, 11], those textual event relations cannot be directly applied to defining the relation among images.
@@
To explore the multi-image perception capabilities of LVLMs, we propose a multi-image relation association task, which requires LVLMs to discern the potential relations between two images (for instance, recognizing that the car and the knife, each present in different images, are both made of iron). We manually curate a Multi-granularity Multi-image Relational Association (MMRA) benchmark, evaluating the multi-image perception capabilities of LVLMs. Based on the relations in ConceptNet and observations of potential connections between images, we define an associational relation system, which consists of 6 subtasks at the entity-level granularity (i.e., RelativePosition, NearSameEntity, MentalitySimilarity, AppearanceSimilarity, SimilarMaterial and UsageSimilarity) and 5 subtasks at the image-level granularity (i.e., Layout, Environments, SimilarEvent, SubEvent and ObservationAngle) across from different perspectives of mining relation between images (outline in Fig 1). Specifically, we select a subset of the images in LLaVA-665k-multi dataset and employ 5 annotators to manually label 1,024 image pairs with questions and answers on the selected data. Besides, to eliminate the answer leakage within the text of the question and its options, we manually remove the content in the question and option, which could make LLMs and VLMs directly infer the answer to the question without analyzing the accompanying images.
To explore how the content captured by visual modules in images affects the multi-image perception capabilities of current LVLMs, we employ the LLaVa-Next-110B model to generate detailed descriptions of the images. We then evaluate both LVLMs and LLMs using our MMRA benchmark across four distinct input configurations: Image+Question (IQ), Description+Question (DQ), Image+Description+Question (IDQ), and Question Only (QO). We present our key insights as follows:
@@
    Based on the results of the IQ and QO setting, we found that closed-source models like GPT-4o, GPT-4v, and Gemini-Flash outperformed all open-source models. In particular, GPT-4o achieved state-of-the-art (SOTA) performance in the overall evaluation. Additionally, different models exhibited significant performance variations across different subtasks of our benchmark. Some open-source models even surpassed GPT-4 in certain subtasks.
@@
    Notably, compared to entity-level tasks, models generally perform better on image-level tasks, and their performance tends to be relatively poor in tasks related to spatial awareness. It indicates that current LVLMs have weak fine-grained multi-image association capabilities and are not proficient in handling spatial perception tasks.
@@
    For image-level tasks, incorporating image descriptions significantly boosts the performance of LLMs, placing them just below GPT-4o and GPT-4v. In contrast, the performance of LVLMs shows no notable improvement with the addition of image descriptions. This indicates that the image-level capability of LVLMs mainly relies on the image content perception ability, and the current LVLMs are limited by the reasoning ability of their language module.
  4. 4.
    Based on our MMRA benchmark, we also examined the multi-image sequence perception capabilities of LVLMs by altering the order of input image pairs. With the exception of Idefics2, most open-source LVLMs scored relatively low overall, suggesting that they inadequately address the modeling of image sequences during the pre-training phase.
These findings highlight the limitations of current LVLMs in comprehending multi-image scenarios. They underscore the complex perceptual demands at the entity level and reveal that the requirements for visual modules diverge when associating multiple images at both the entity and image levels.
@@ 2.1 Image Pair Selection
Given that most tasks in the MMRA benchmark require a specific relationship between paired images, we use the semantic similarity of image captions to identify and select image pairs with relatively higher relevance. This aims to reduce the complexity of annotation. To be specific, we randomly chose the images in the LLaVA-665k-multi to form an image pair. We then utilize the SentenceBERT[19] to calculate the semantic similarity to filter the image pair with a score below 0.5. Finally, we obtained 3,403 image pairs for annotation.
@@ 2.2 Subtask Definition
As shown in the Fig 7 of Appendix C, based on the perspective of humans observing images, we divide our tasks into two granularity levels (i.e., entity and the whole image). Because the ConceptNet comprehensively defines the relations among different textual event and entity, most of our subtasks, both at entity level and image level, are extended from the relations in ConceptNet. Additionally, we have also designed some subtasks from a visual perspective (i.e., Layout and ObservationAngle).
Entity level.
@@
  • NearSameEntity (NSE): The relation ’LocatedNear’ in ConceptNet expresses “A and B are typically found near each other”. Based on it, we design a subtask, ’NearSameEntity’, which requires LVLMs to determine whether there are entities, respectively in different images, near the same object.
  • MentalitySimilarity (MS): ’HasProperty’ in ConceptNet is a relation that describes the characteristics of an entity. We think the emotional property expressed by the images could directly affect humans. Thus, we extend this relation to a subtask that requires LVLMs to determine whether the creatures in two images have similar emotions, attitudes, or feelings (e.g., happy, excited, serious, surprised, etc.).
@@
  • SimilarMaterial (SM): The relation ’MadeOf’ in ConceptNet expresses ’A is made of B’. Therefore, we design the subtask ’SimilarMaterial’ which requires LVLMs to judge whether there are entities, respectively in different images, with the same production materials.
@@
  • Layout (LO): At the image granularity, we regard the layout of the image as a representation of the relation “AtLocation”. We design a subtask that requires the LVLMs to determine whether there are similarities in layout between images according to the relation ’NearBy’.
  • Environment (Env): From the visual perspective, the environment of the image is also an important content that humans tend to notice (e.g., both images depict the streets of a European country with a Gothic architectural style). So, we design a subtask that lets LVLMs judge if the environments in those images are similar according to the relation ’AtLocation’.
  • SimilarEvent (SimE): Excepting the ’SubEvent’ relation in ConceptNet, the similar event is also a crucial factor. So we devise a subtask to evaluate the LVLMs’ capability to find the same event that happened in the given two images.
  • SubEvent (SubE): The temporary relation is an important connection between two images. Therefore, we extend the relation ’SubEvent’ to a subtask that requires LVLMs to determine whether the two images describe events that occurred at the same scene in two consecutive moments.
@@ 2.3 Data Annotation
As shown in Fig 2, each annotator is provided with two images and a certain subtask (i.e., Environment). Their responsibility is to determine whether they could design a question based on the given task for the image pair. If the image pair meets the task requirements, they proceed to annotate a question, and options (either multiple-choice or true/false) for that pair. The annotator terminates annotating a task once they reach a predetermined number of labelled samples (i.e., 90) or once all the image pairs for that task have been annotated.
Figure 2: The process of annotation.
@@
We also conducted cross-validation on the annotated data. Specifically, each annotator reviews 2-3 tasks labeled by their peers. If any annotated samples don’t meet the task requirements or if the answers derived from the images and options don’t match the correct answer, those samples are removed. Additionally, we ensure that the answers to all questions cannot be directly inferred from the text alone (i.e., the question and options). Quality control is concluded once all annotators agree that their verified portion satisfies the specified requirements.
@@
As shown in Fig 3, we obtain a total of 1,024 annotated samples. To maintain the balance of samples of the subtasks, we endeavored to maintain that the number of samples for all tasks is around 90. The ObservationAngle task has the highest proportion in the entire benchmark, with a total of 126 samples (12.28%). Due to the difficulty of labeling in the NearSameEntity task, we removed some samples with inconsistent opinions from different annotators during the quality control process and this subtask only has 65 samples.
@@ 2.4 Elimination of Answer Leakage from Questions and Options
When designing multiple-choice options at the entity level, we need to identify potential entities that could be regarded as the correct answer to the question and provide justifications. For example, as illustrated in Fig 1, ’both tree and bench are made of wood’ can be the answer to the SimilarMaterial subtask. However, language models can sometimes deduce the correct answer simply by analyzing the textual content in options. Additionally, annotators often unconsciously label the correct answer with greater detail and specificity, and the language model towards choosing these more detailed options.
@@
Figure 4: Comparing results before and after textual answer leakage elimination.
@@ 3.1 Experiments Setting
To explore the impact of LVLM’s image-captioning ability on its multi-image perception, we design four input settings: Image + Question, Description + Question, Image + Description + Question and Question Only. Specifically, we use the LLaVA-NeXT-100B to obtain a detailed textual description of the image in the MMRA benchmark.
@@
For the sake of inspecting whether the answer to the questions in our benchmark is leaked in the textual information of options and questions, we only input the question to let LVLMs answer.
@@ 3.2 Baselines
As shown in Tab 1, we evaluated our benchmark on both mainstream closed-source and open-source large models. Regarding close-source LVLMs, we choose OpenAI’s GPT4o and GPT4v, as well as Google’s Gemini-Flash. As for the open-source LVLMs, we mainly evaluate those supporting multi-image inputs (i.e., Idefics2 , Qwen-VL-Chat, Phi3v, Mantis-Idefics2). Besides, we also assess the open-source LLMs (i.e., LLaMA, Qwen1.5-Chat, Qwen2-Chat) under the text-only input setting.
@@ 3.3 Evaluation Protocol
As for each task, we all design a prompt to make LVLMs directly generate textual format answers to the question. Except including the content described in the section 3.1, we let LVLMs generate the ’A’, ’B’, ’C’ or ’D’ for the choice questions, and ’Yes’ or ’No’ for the T/F questions. Besides, we also add the options to the prompt. As for further details about our prompt design, please refer to the Tab 5 in Appendix B.
@@
Because the golden answer in our benchmark is in the format of option id (i.e., ’A’, ’B’, ’C’ and ’D’) or judgment (i.e., ’Yes’ or ’No’), we design a rule to match the response of LVLMs with the golden answer. Finally, we use the accuracy of the matching results as the score of those models. Please refer to the Appendix A for the details of our designed matching rule.